# Learning to Segment for Vehicle Routing Problems

### v1

@@ 5.1 L2Seg Accelerates Various Iterative Backbone Solvers
We first verify the effectiveness of the three proposed L2Seg variants to enhance backbone iterative solvers. Table 1 presents results on large-capacity, uniformly distributed CVRPs with long subtours. Results show that all L2Seg variants consistently improve each backbone solver across all problem scales. We also note that performance gains are larger for weaker backbones, and while L2Seg-AR and L2Seg-NAR each boost performance, their combination (L2Seg-SYN) delivers the best solutions. Figure 4 plots average objective curves over time, which reveals 2x to 7x speedups on the backbone iterative solvers with L2Seg-SYN. Remarkably, L2Seg-augmentation lets weaker solvers surpass stronger ones (e.g., LKH-3 + L2Seg-SYN outperforms vanilla LNS).
@@
HGS [8] 41.20 0.00% 5m 57.20 0.00% 5m 126.20 0.00% 5m
LKH-3 [7] 42.98 4.32% 6.6m 57.94 1.29% 11.4m 175.70 39.22% 2.5m
LNS [9] 42.44 3.01% 2.5m 57.62 0.73% 4.0m 126.58 0.30% 5.0m
BQ [16] 44.17 7.21% 55s 62.59 9.42% 3m 139.80 10.78% 45m
LEHD [25] 43.96 6.70% 1.3m 61.58 7.66% 9.5m 138.20 9.51% 3h
ELG [15] 43.58 5.78% 15.6m - - - - - -
ICAM [51] 43.07 4.54% 26s 61.34 7.24% 3.7m 136.90 8.48% 50m
L2R [42] 44.20 7.28% 34.2s - - - 131.10 3.88% 1.8m
SIL [5] 42.00 1.94% 1.3m 57.10 -0.17% 2.4m 123.10 -2.52% 5.9m
TAM(LKH-3) [41] 46.30 12.38% 4m 64.80 13.29% 9.6m 144.60 14.58% 35m
GLOP-G(LKH-3) [40] 45.90 11.41% 2m 63.02 10.52% 2.5m 140.40 11.25% 8m
UDC [26] 43.00 4.37% 1.2h 60.01 4.9% 2.15h 136.70 8.32% 16m
L2D [4] 42.07 2.11% 2.5m 57.44 0.42% 4.2m 126.48 0.22% 5.3m
@@
5.2 L2Seg Outperforms Both Classic and Neural Baselines on CVRP and VRPTW
We evaluate the highest-performing L2Seg-SYN implementation with three distinct backbone solvers and compare against state-of-the-art classical and neural approaches. As demonstrated in Table 2, L2Seg surpasses both classical and neural baselines on CVRP and VRPTW benchmarks. For CVRP, L2Seg achieves superior performance within comparable computational time relative to competitive classical solvers, including HGS on larger problem instances. It also outperforms the state-of-the-art learning-based constructive solver SIL [5] and divide-and-conquer solver L2D [4] across all problem scales. For VRPTW, L2Seg exceeds all classical and learning-based solvers by at least $3\%$ across various scales under identical time constraints, with performance advantages increasing as problem size grows. Notably, L2Seg consistently enhances performance when integrated with any backbone solver, demonstrating its versatility. Additional analyses are provided in Appendix C.
Table 3: Performance of L2Seg-SYN v.s. Random FSTA to accelerate LNS on CVRP instances.
@@ 5.3 Further Analysis and Discussions
Ablation Study. Table 3 compares the LNS backbone; random FSTA with 40% and 60% of edges arbitrarily marked as unstable; L2Seg-SYN w/o enhanced features; and full L2Seg-SYN. Results show that Random FSTA worsens performance; and only full L2Seg-SYN with enhanced features achieves top performance. This confirms that L2Seg’s learnable, feature-guided segmentation is indispensable for preserving high-quality segments in FSTA for boosting backbone solvers.
Figure 6: Problem size reduction and prediction statistics across L2Seg configurations (different problem reduction strategies). Size represents the ratio of reduced to original problem dimensions.
Why NAR+AR Is The Best Performance? We offer the following insights: L2Seg-NAR effectively identifies unstable regions but lacks the ability to model dependencies among unstable edges, often over-classifying within these regions. L2Seg-AR captures local dependencies but struggles to detect unstable regions initially. L2Seg-SYN combines these strengths, identifying unstable regions through its NAR component and refining predictions via its AR component. We provide a conceptual illustration in Figure 5, with actual instance visualizations in Appendix D. Table 4 presents results on CVRP2k, where 16.4% of edges are unstable. Higher Recall enables more edges to be reoptimized, potentially improving solution quality, while higher TNR reduces problem size and runtime. L2Seg-NAR yields higher Recall but lower TNR, leading to longer runtimes. L2Seg-AR underperforms in both metrics, confirming its difficulty in identifying unstable regions.
Pursuing High Recall v.s. High TNR. Higher Recalls allow more unstable edges to be reoptimized, potentially improving solution quality, while higher TNRs reduce problem size and runtime. However, due to imprecision in learning, pursuing high TNRs often reduces Recall, causing premature convergence to local optima. Figure 6(b) shows that fixing too few edges (left: high Recall, low TNR) or too many (right: high TNR, low Recall) degrades performance. Our configuration (middle) balances this trade-off for optimal performance.
@@ 6 Conclusion
This work introduces Learning-to-Segment (L2Seg), a novel learning-guided framework that accelerates state-of-the-art iterative solvers for large-scale VRPs by up to 7x. We first formalize the FSTA decomposition and then employ a specialized encoder-decoder architecture to dynamically differentiate potentially unstable and stable segments in FSTA. L2Seg features three model variants, L2Seg-NAR, L2Seg-AR, and L2Seg-SYN, pioneering the synergy of AR and NAR models in NCO. Extensive results show that L2Seg demonstrates state-of-the-art performance on representative CVRP and VRPTW and also exhibits flexibility in boosting traditional and learning-based solvers, including other decomposition frameworks. One potential limitation is that L2Seg is not guaranteed to boost all VRP solvers across all VRP variants. Future work includes: (1) extending L2Seg to accelerate a broader range of VRP solvers (e.g., [8, 53]); (2) applying L2Seg to additional VRP variants and other combinatorial optimization problems; (3) expanding the synergy between AR and NAR models to the broader NCO community; (4) distilling learned behavior via Large Language Models (LLMs) [54]; and (5) exploring the explainability and theoretical foundations of the learned behavior.
Acknowledgments and Disclosure of Funding
@@
Appendix A Details of First-Segment-Then-Aggregate (FSTA)
A.1 More discussions on FSTA
A.1.1 Visualization of Unstable Edge Patterns
@@
A.1.2 Visualization of Applying FSTA on One CVRP Instance
@@
Figure 8: Illustration of our FSTA applied to one CVRP instance. Each FSTA step corresponds to the descriptions in Appendix A.1.4. Red dashed lines: unstable edges; blue dashed lines: re-optimized edges. Note that the subproblem (d) contains substantially fewer nodes than the original instance (a).
To provide a concrete illustration of our FSTA methodology, we present an example of its application to CVRP in Figure 8, which demonstrates the complete FSTA decomposition pipeline (detailed algorithmic specifications are provided in Appendix A.1.4). This example utilizes the lookahead oracle model for unstable edge identification (defined in Appendix A.1.1), employs LKH-3 as the backbone optimization solver, and operates on a representative small-capacity CVRP1k instance to showcase the framework’s efficacy. Red dashed lines indicate detected unstable edges, while blue dashed lines represent re-optimized edges. Note that dual hypernode aggregation substantially reduces the problem size compared to the original instance.
A.1.3 Assumption Verification
@@
A.1.4 Details of FSTA Decomposition Framework
@@
    Unstable Edges Detection: We implement effective methods (e.g., our learning-based model L2Seg or random heuristics detailed in Section 5.3) to identify unstable edges $E_{\text{unstable}}$ and obtain the stable edge set $E_{\text{stable}}=E\setminus E_{\text{unstable}}$. This identification challenge is addressed by our L2Seg model, with full details provided in Section 4 and Appendix B.
@@
    Re-optimization with Backbone Solvers: We invoke a backbone solver to improve solution $\tilde{\mathcal{R}}$, yielding an enhanced solution $\tilde{\mathcal{R}}_{+}$. While theoretically any solver could serve as the backbone solver, practical acceleration requires solvers capable of effectively leveraging existing solutions (e.g., LKH-3 [7]).
@@
Selection of Backbone Solvers. Our framework is generic to be applied to most existing VRP heuristics by design. In practice, acceleration within our framework requires solvers that can effectively utilize initial solutions as warm starts. Furthermore, if the dual hypernode aggregation is used, the backbone solver needs to fix certain edges during local search. Our framework is readily compatible with a variety of solvers without modifying their source codes, including LKH-3 [7], decomposition-based solvers like LNS [9], and learning-based methods such as L2D [4]. Incorporating additional solvers such as HGS [8], would involve extending its current code to accept initial solutions as input, which we leave as future work. Notably, as demonstrated in Section 5, our L2Seg-augmented approach with relatively weaker backbone solvers outperforms HGS in multiple CVRP and VRPTW benchmark scenarios.
Applicability to Routing Variants. FSTA is broadly applicable to routing problem variants that support feasible hypernode aggregation and solution recovery, as ensured by the feasibility and monotonicity conditions established in Section 3. In Appendix A.2, we formally prove that many routing variants meet these conditions, demonstrating the versatility of our L2Seg framework. Detailed implementation guidelines for applying hypernode aggregation across different routing variants are provided in Appendix A.1.5.
@@ 5       $\{S^{i}_{j,k}\}\leftarrow\text{GetSegments}(P,\mathcal{R},E_{\text{unstable}})$ ;
9       ~$\tilde{\mathcal{R}}_{+}\leftarrow\text{BS}(\tilde{P},\tilde{\mathcal{R}})$ ;
@@
11       $\mathcal{R}_{+}\leftarrow\text{RecoverSolution}(P,\tilde{P},\tilde{\mathcal{R}}_{+})$ ;
@@ 13       $\mathcal{R}\leftarrow\mathcal{R}_{+}$ ;
A.1.5 Applying FSTA on Various VRPs
In this section, we present the implementation details of FSTA across diverse routing variants, including the Capacitated Vehicle Routing Problem (CVRP), Vehicle Routing Problem with Time Windows (VRPTW), Vehicle Routing Problem with Backhauls (VRPB), Single-Commodity Vehicle Routing Problem with Pickup and Delivery (1-VRPPD), and Risk-Constrained Cash-in-Transit Vehicle Routing Problem (RCTVRP). Without loss of generality, we denote a segment to be aggregated as $S_{j,k}=(x_{j}\rightarrow\ldots\rightarrow x_{k})$, and its corresponding hypernode representation as either $\tilde{S}_{j,k}=\{\tilde{x}\}$ (single hypernode) or $\tilde{S}_{j,k}=\{\tilde{x}_{j},\tilde{x}_{k}\}$ (dual hypernodes). The implementation specifications are summarized in Table 6.
@@ A.1.5 Applying FSTA on Various VRPs
VRPTW. We provide the formal definition of VRPTW in Section 3. In addition to location and demand attributes, VRPTW instances are characterized by time windows $[t^{l},t^{r}]$ and service time $s$ for each node. For VRPTW, we employ adaptive strategies for hypernode aggregation based on temporal feasibility. We first compute the aggregated time windows $\bar{t}^{l}_{j}$, $\bar{t}^{r}_{j}$ and aggregated service time $\bar{s}_{j}$ using the following recursive formulation:
$\bar{t}_{m}^{l}$ =$=\begin{cases}t_{k}^{l}&\text{if }m=k\\ \max\{t_{m}^{l},\bar{t}_{m+1}^{l}-(s_{m}+\text{dist}(x_{m},x_{m+1}))\}&\text{if }j\leq m\leq k-1,\end{cases}$ (3)
$\bar{t}_{m}^{r}$ =$=\begin{cases}t_{k}^{r}&\text{if }m=k\\ \min\{t_{m}^{r},\bar{t}_{m+1}^{r}-(s_{m}+\text{dist}(x_{m},x_{m+1}))\}&\text{if }j\leq m\leq k-1,\end{cases}$
$\bar{s}_{m}$ =$=\begin{cases}s_{k}&\text{if }m=k\\ \bar{s}_{m+1}+s_{m}+\text{dist}(x_{m},x_{m+1})&\text{if }j\leq m\leq k-1,\end{cases}$
where $[t_{m}^{l},t_{m}^{r}]$ denotes the time window for node $x_{m}$, $s_{m}$ represents the service time at node $x_{m}$, and $\text{dist}(x_{m},x_{m+1})$ is the travel time from node $x_{m}$ to node $x_{m+1}$.
@@
VRPB. Compared to the CVRP, The VRPB [55] involves serving two types of customers: linehaul customers requiring deliveries from the depot and backhaul customers providing goods to be collected and returned to the depot. The primary constraint is that all linehaul customers must be visited before any backhaul customers on the same route, while ensuring vehicle capacity is never exceeded during either the delivery or pickup phases. We use $b_{i}\in\{0,1\}$ to indicate whether node $i$ is a backhaul customer. For VRPB, we require the edge connecting to a linehaul customer and a backhaul customer included in the $E_{\text{unstable}}$. We employ single hypernode aggregation that $\text{dist}(x_{i},\tilde{x})=\text{dist}(x_{i},x_{j})$, $\text{dist}(\tilde{x},x_{i})=\text{dist}(x_{k},x_{i})$, $\tilde{d}=d_{j}+\cdots+d_{k}$, and $\tilde{b}=b_{j}$ (we require customer being the same type within each segment that $b_{j}=...=b_{k}$).
Table 6: Implementation specifications of FSTA hypernode aggregation for CVRP, VRPTW, and VRPB variants. Refer to Equation 3 for the definitions of $\bar{s}_{j}$, $\bar{t}^{l}_{j}$ and $\bar{t}^{r}_{j}$.
@@
A.2 Proof of FSTA
Theorem. (Feasibility) If the aggregated solution $\tilde{\mathcal{R}}_{+}$ is a feasible solution to the aggregated problem, then $\mathcal{R}_{+}$ is a feasible solution to the original, non-aggregated problem. (Monotonicity) Let ~+$\tilde{\mathcal{R}}^{1}_{+}$ and ~+$\tilde{\mathcal{R}}^{2}_{+}$ be two feasible solutions to the aggregated problem, with $f(\tilde{\mathcal{R}}^{1}_{+})\leq f(\tilde{\mathcal{R}}^{2}_{+})$, where $f(\cdot)$ denotes the objective function (total travel cost). Then, for the associated solution in the original space, we also have $f(\mathcal{R}^{1}_{+})\leq f(\mathcal{R}^{2}_{+})$.

### v2

@@ 5.1 L2Seg Accelerates Various Iterative Backbone Solvers
We first verify the effectiveness of the three L2Seg variants to enhance backbone solvers. Table 1 presents results on large‑capacity, uniformly distributed CVRPs with long subtours. All L2Seg variants consistently improve each backbone across all problem scales. Also, performance gains are larger for weaker backbones. While L2Seg‑AR and L2Seg‑NAR each boost performance, their combination (L2Seg‑SYN) delivers the best solutions. Figure 5 plots average objective curves over time, which reveal 2x to 7x speedups on the backbone solvers with L2Seg‑SYN. Remarkably, L2Seg‑augmentation lets weaker solvers surpass stronger ones (e.g., LKH‑3 + L2Seg‑SYN outperforms vanilla LNS).
@@
HGS (Vidal, 2022) 41.20 0.00% 5m 57.20 0.00% 5m 126.20 0.00% 5m
LKH-3 (Helsgaun, 2017) 42.98 4.32% 6.6m 57.94 1.29% 11.4m 175.70 39.22% 2.5m
LNS (Shaw, 1998) 42.44 3.01% 2.5m 57.62 0.73% 4.0m 126.58 0.30% 5.0m
BQ (Drakulic et al., 2023) 44.17 7.21% 55s 62.59 9.42% 3m 139.80 10.78% 45m
LEHD (Luo et al., 2023) 43.96 6.70% 1.3m 61.58 7.66% 9.5m 138.20 9.51% 3h
ELG (Gao et al., 2024) 43.58 5.78% 15.6m - - - - - -
ICAM (Zhou et al., 2024) 43.07 4.54% 26s 61.34 7.24% 3.7m 136.90 8.48% 50m
L2R (Zhou et al., 2025a) 44.20 7.28% 34.2s - - - 131.10 3.88% 1.8m
SIL (Luo et al., 2024) 42.00 1.94% 1.3m 57.10 -0.17% 2.4m 123.10 -2.52% 5.9m
TAM(LKH-3) (Hou et al., 2023) 46.30 12.38% 4m 64.80 13.29% 9.6m 144.60 14.58% 35m
GLOP-G(LKH-3) (Ye et al., 2024) 45.90 11.41% 2m 63.02 10.52% 2.5m 140.40 11.25% 8m
UDC (Zheng et al., 2024) 43.00 4.37% 1.2h 60.01 4.9% 2.15h 136.70 8.32% 16m
L2D (Li et al., 2021) 42.07 2.11% 2.5m 57.44 0.42% 4.2m 126.48 0.22% 5.3m
NDS (Hottung et al., 2025) 41.16 -0.01% 2.5m 56.11 -1.91% 4m - - -
@@
5.2 L2Seg Outperforms Classic and Neural Baselines on CVRP and VRPTW
We evaluate the highest-performing L2Seg-SYN implementation with three distinct backbone solvers and compare against state-of-the-art classical and neural approaches. As demonstrated in Table 2, L2Seg surpasses both classical and neural baselines on CVRP and VRPTW benchmarks. For CVRP, L2Seg achieves superior performance within comparable computational time relative to competitive classical solvers, including HGS on larger problem instances. It also outperforms the state-of-the-art learning-based constructive solver SIL (Luo et al., 2024) and divide-and-conquer solver L2D (Li et al., 2021) across all problem scales. For VRPTW, L2Seg exceeds all classical and learning-based solvers across various scales under identical time constraints, with performance advantages increasing as problem size grows. Notably, L2Seg consistently enhances performance when integrated with any backbone solver, demonstrating its versatility. Additional analyses are provided in Appendix E.
Table 3: Performance of L2Seg‑SYN v.s. Random FSTA to accelerate LNS on CVRP instances.
@@ 5.3 Further Analysis and Discussions
Ablation Study. Table 3 compares the LNS backbone; random FSTA with 40% and 60% of edges arbitrarily marked as unstable; L2Seg‑SYN w/o enhanced features; and full L2Seg‑SYN. Results show that Random FSTA worsens performance; and only full L2Seg‑SYN with enhanced features achieves top performance. This confirms that L2Seg’s learnable, feature‑guided segmentation is indispensable for preserving high‑quality segments in FSTA for boosting backbone solvers.
Figure 7: Statistic values of Size (reduced/original ratio), Recall, and TNR across three L2Seg-SYN configurations.
High Recall or High TNR? Higher Recall allows more unstable edges to be reoptimized, potentially improving performance, while higher TNR reduces problem size and runtime. However, due to learning imprecision, pursuing high TNRs often reduces Recall, causing premature convergence. Figure 7 shows that for L2Seg-SYN, fixing too few (left: high Recall, low TNR) or too many (right: high TNR, low Recall) degrades performance. Ours (middle) balances this tradeoff for optimal performance.
Why NAR+AR Is the Best? Figure 6 shows a conceptual illustration of the model’s behaviour across L2Seg variants (See Appendix E.5 for a real case-study). L2Seg-NAR identifies unstable regions but over-classifies due to the lack of dependency modeling, while L2Seg-AR models dependencies but struggles with initial detection. L2Seg-SYN achieves the complementary synergy. Moreover, Table 4 further shows that L2Seg-SYN achieves the best balanced Recall and TNR for the best performance.
@@ 6 Conclusion
This work introduces Learning-to-Segment (L2Seg), a novel learning-guided framework that accelerates state-of-the-art iterative solvers for large-scale VRPs by 2x to 7x. We formalize the FSTA decomposition and employ a specialized encoder-decoder architecture to dynamically differentiate potentially unstable and stable segments in FSTA. L2Seg features three variants, L2Seg-NAR, L2Seg-AR, and L2Seg-SYN, pioneering the synergy of AR and NAR models in NCO. Extensive results demonstrate L2Seg’s state-of-the-art performance on representative CVRP and VRPTW and flexibility in boosting classic and learning-based solvers, including other decomposition frameworks. One potential limitation is that L2Seg is not guaranteed to boost all VRP solvers across all VRP variants. Future work includes: (1) extending L2Seg to accelerate additional VRP solvers (e.g., Vidal (2022)); (2) applying L2Seg to more VRP variants and other combinatorial optimization problems; and (3) expanding the synergy between AR and NAR models to the broader NCO community.
Reproducibility Statement
We provide comprehensive technical details in the appendices: architecture and input features (Appendix D.3), data generation (Appendix D.4), training procedures (Appendix C.4), and experimental setup (Section 5). The complete codebase, including code and pre-trained models, will be released on GitHub under the MIT License upon publication.
Acknowledgment
@@
Appendices
Contents
Appendix A Supplementary Definitions
A.1 Unstable Edges and Stable Edges
Unstable edges refer to edges that need to be re-optimized during the iterative re-optimization procedure. We supplement the formal definitions as follows: given a solution $\mathcal{R}_{t}$ at iterative step $t$, an edge $e\in\mathcal{R}$ is unstable if $e\notin\mathcal{R}_{t+1}$ or $e\notin\mathcal{R}_{t+2}$, …, or $e\notin\mathcal{R}_{t+k}$. When we generate the labels for training, we use a lookahead backbone solver for detecting unstable edges, which equivalently sets $k=1$. An edge is a stable edge if it’s not an unstable edge.
A.2 Capacitated Vehicle Routing Problem
Given a complete graph $G=(V,E)$ where $V=\{x_{0},x_{1},\ldots,x_{n}\}$ is the set of nodes with node $x_{0}$ representing the depot and nodes $x_{1}$ to $x_{n}$ representing customers. Each customer $i$ has a demand $d_{i}>0$, and each edge $e_{i,j}\in E$ has an associated cost representing the travel distance or travel time between nodes $x_{i}$ and $x_{j}$. A fleet of homogeneous vehicles, each with capacity $C$, is available at the depot. The objective is to find a set of routes that minimizes the total travel cost, subject to: (i) each route starts and ends at the depot, (ii) each customer is visited exactly once, (iii) the total demand of customers on each route does not exceed vehicle capacity $C$.
A.3 Vehicle Routing Problem with Time Windows
Given a complete graph $G=(V,E)$ where $V=\{x_{0},x_{1},\ldots,x_{n}\}$ is the set of nodes with node $x_{0}$ representing the depot and nodes $x_{1}$ to $x_{n}$ representing customers. Each customer $i$ has a demand $d_{i}>0$, and each edge $e_{i,j}\in E$ has an associated cost representing the travel distance or travel time between nodes $x_{i}$ and $x_{j}$. Each customer $i$ has a time window $t^{l}_{i},t^{r}_{i}]$ where $t^{l}_{i}$ is the earliest arrival time and $t^{r}_{i}$ is the latest arrival time, and requires a service time $s_{i}$. A fleet of homogeneous vehicles, each with capacity $C$, is available at the depot. The objective is to find a set of routes that minimizes the total travel cost, subject to: (i) each route starts and ends at the depot, (ii) each customer is visited exactly once, (iii) the total demand of customers on each route does not exceed vehicle capacity $C$, (iv) service at each customer begins within their time window $t^{l}_{i},t^{r}_{i}]$.
Appendix B Details of First-Segment-Then-Aggregate (FSTA)
B.1 More discussions on FSTA
B.1.1 Visualization of Unstable Edge Patterns
@@
B.1.2 Visualization of Applying FSTA on One CVRP Instance
@@
Figure 9: Illustration of our FSTA applied to one CVRP instance. Each FSTA step corresponds to the descriptions in Appendix B.1.4. Red dashed lines: unstable edges; blue dashed lines: re-optimized edges. Note that the subproblem (d) contains substantially fewer nodes than the original instance (a).
To provide a concrete illustration of our FSTA methodology, we present an example of its application to CVRP in Figure 9, which demonstrates the complete FSTA decomposition pipeline (detailed algorithmic specifications are provided in Appendix B.1.4). This example utilizes the lookahead oracle model for unstable edge identification (defined in Appendix B.1.1), employs LKH-3 as the backbone optimization solver, and operates on a representative small-capacity CVRP1k instance to showcase the framework’s efficacy. Red dashed lines indicate detected unstable edges, while blue dashed lines represent re-optimized edges. Note that dual hypernode aggregation substantially reduces the problem size compared to the original instance.
B.1.3 Assumption Verification
@@
B.1.4 Details of FSTA Decomposition Framework
@@
    Unstable Edges Detection: We implement effective methods (e.g., our learning-based model L2Seg or random heuristics detailed in Section 5.3) to identify unstable edges $E_{\text{unstable}}$ and obtain the stable edge set $E_{\text{stable}}=E\setminus E_{\text{unstable}}$. This identification challenge is addressed by our L2Seg model, with full details provided in Section 4 and Appendix C.
@@
    Re-optimization with Backbone Solvers: We invoke a backbone solver to improve solution $\tilde{\mathcal{R}}$, yielding an enhanced solution $\tilde{\mathcal{R}}_{+}$. While theoretically any solver could serve as the backbone solver, practical acceleration requires solvers capable of effectively leveraging existing solutions (e.g., LKH-3 (Helsgaun, 2017)).
@@
Selection of Backbone Solvers. Our framework is generic to be applied to most existing VRP heuristics by design. In practice, acceleration within our framework requires solvers that can effectively utilize initial solutions as warm starts. Furthermore, if the dual hypernode aggregation is used, the backbone solver needs to fix certain edges during local search. Our framework is readily compatible with a variety of solvers without modifying their source codes, including LKH-3 (Helsgaun, 2017), decomposition-based solvers like LNS (Shaw, 1998), and learning-based methods such as L2D (Li et al., 2021). Incorporating additional solvers such as HGS (Vidal, 2022), would involve extending its current code to accept initial solutions as input, which we leave as future work. Notably, as demonstrated in Section 5, our L2Seg-augmented approach with relatively weaker backbone solvers outperforms HGS in multiple CVRP and VRPTW benchmark scenarios.
Applicability to Routing Variants. FSTA is broadly applicable to routing problem variants that support feasible hypernode aggregation and solution recovery, as ensured by the feasibility and monotonicity conditions established in Section 3. In Appendix B.2, we formally prove that many routing variants meet these conditions, demonstrating the versatility of our L2Seg framework. Detailed implementation guidelines for applying hypernode aggregation across different routing variants are provided in Appendix B.1.5.
@@ 5 $\{S^{i}_{j,k}\}\leftarrow\text{GetSegments}(P,\mathcal{R},E_{\text{unstable}})$ ;
[…]
@@
[…]
@@ 13 $\mathcal{R}\leftarrow\mathcal{R}_{+}$ ;
B.1.5 Applying FSTA on Various VRPs
In this section, we present the implementation details of FSTA across diverse routing variants, including the Capacitated Vehicle Routing Problem (CVRP), Vehicle Routing Problem with Time Windows (VRPTW), Vehicle Routing Problem with Backhauls (VRPB), and Single-Commodity Vehicle Routing Problem with Pickup and Delivery (1-VRPPD). Without loss of generality, we denote a segment to be aggregated as $S_{j,k}=(x_{j}\rightarrow\ldots\rightarrow x_{k})$, and its corresponding hypernode representation as either $\tilde{S}_{j,k}=\{\tilde{x}\}$ (single hypernode) or $\tilde{S}_{j,k}=\{\tilde{x}_{j},\tilde{x}_{k}\}$ (dual hypernodes). The implementation specifications are summarized in Table 6.
@@ B.1.5 Applying FSTA on Various VRPs
VRPTW. We provide the formal definition of VRPTW in Section 3. In addition to location and demand attributes, VRPTW instances are characterized by time windows $t^{l},t^{r}]$ and service time $s$ for each node. For VRPTW, we employ adaptive strategies for hypernode aggregation based on temporal feasibility. We first compute the aggregated time windows $\bar{t}^{l}_{j}$, $\bar{t}^{r}_{j}$ and aggregated service time $\bar{s}_{j}$ using the following recursive formulation:
[…]
where $t_{m}^{l},t_{m}^{r}]$ denotes the time window for node $x_{m}$, $s_{m}$ represents the service time at node $x_{m}$, and $\text{dist}(x_{m},x_{m+1})$ is the travel time from node $x_{m}$ to node $x_{m+1}$.
@@
VRPB. Compared to the CVRP, the VRPB (Goetschalckx and Jacobs-Blecha, 1989) involves serving two types of customers: linehaul customers requiring deliveries from the depot and backhaul customers providing goods to be collected and returned to the depot. The primary constraint is that all linehaul customers must be visited before any backhaul customers on the same route, while ensuring vehicle capacity is never exceeded during either the delivery or pickup phases. We use $b_{i}\in\{0,1\}$ to indicate whether node $i$ is a backhaul customer. For VRPB, we require the edge connecting to a linehaul customer and a backhaul customer included in the $E_{\text{unstable}}$. We employ single hypernode aggregation that $\text{dist}(x_{i},\tilde{x})=\text{dist}(x_{i},x_{j})$, $\text{dist}(\tilde{x},x_{i})=\text{dist}(x_{k},x_{i})$, $\tilde{d}=d_{j}+\cdots+d_{k}$, and $\tilde{b}=b_{j}$ (we require customer being the same type within each segment that $b_{j}=...=b_{k}$).
1-VRPPD. Compared to the CVRP, the 1-VRPPD (Martinovic et al., 2008) deals with customers labeled as either cargo sink ($d_{i}<0$) or cargo source ($d_{i}>0$), depending on their pickup or delivery demand. Along the route of each vehicle, the vehicle could not load negative cargo or cargo exceeding the capacity of the vehicle $C$. For any segment $S_{j,k}$, we define $D^{j}=d_{j}$, $D^{j+1}=d_{j}+d_{j+1}$, …, and $D^{k}=d_{j}+d_{j+1}+...+d_{k}$. We further define $D^{\text{min}}=\min\{0,D_{j},D_{j+1},$...\} and $D^{\text{max}}=\max\{0,D_{j},D_{j+1},$...\}. For 1-VRPPD, we require three hypernodes $\tilde{x}_{j}=x_{j}$, $\tilde{x}_{\text{mid}}$, and $\tilde{x}_{k}=x_{k}$, where the distances from $\tilde{x}_{\text{mid}}$ to $\tilde{x}_{j}$ or $\tilde{x}_{k}$ are 0, and infinity for the other hypernodes. For the aggregated demands, $\tilde{d}_{j}=D^{\text{min}}$, $\tilde{d}_{\text{mid}}=D^{\text{max}}-D^{\text{min}}$, and $\tilde{d}_{k}=D^{k}-D^{\text{max}}-D^{\text{min}}$. Additional constraints are added to ensure the directed edges $\tilde{x}_{j}\rightarrow\tilde{x}_{\text{mid}}\rightarrow\tilde{x}_{k}$ are included in the solutions.
Table 6: Implementation specifications of FSTA hypernode aggregation for CVRP, VRPTW, VRPB variants. Refer to Equation 3 for the definitions of $\bar{s}_{j}$, $\bar{t}^{l}_{j}$ and $\bar{t}^{r}_{j}$.
@@
1-VRPPD
Three Hypernodes Always Location/Distance $\tilde{x}_{j}=x_{j}$, $\tilde{x}_{k}=x_{k}$ Include edges $\tilde{x}_{j}\rightarrow\tilde{x}_{\text{mid}}\rightarrow\tilde{x}_{k}$ in the solution
$\text{dist}(\tilde{x}_{j},\tilde{x}_{\text{mid}})=\text{dist}(\tilde{x}_{\text{mid}},\tilde{x}_{k})=0$
$\tilde{x}_{\text{mid}}$ only connects to $\tilde{x}_{j}$ and $\tilde{x}_{k}$
Demand $\tilde{d}_{j}=D^{\text{min}}$, $\tilde{d}_{\text{mid}}=D^{\text{max}}-D^{\text{min}}$,
$\tilde{d}_{k}=D^{k}-D^{\text{max}}-D^{\text{min}}$
B.2 Proof of FSTA
Theorem. (Feasibility) If the aggregated solution $\tilde{\mathcal{R}}_{+}$ is a feasible solution to the aggregated problem, then $\mathcal{R}_{+}$ is a feasible solution to the original, non-aggregated problem. (Monotonicity) Let $\tilde{\mathcal{R}}^{1}_{+}$ and $\tilde{\mathcal{R}}^{2}_{+}$ be two feasible solutions to the aggregated problem, with $f(\tilde{\mathcal{R}}^{1}_{+})\leq f(\tilde{\mathcal{R}}^{2}_{+})$, where $f(\cdot)$ denotes the objective function (total travel cost). Then, for the associated solution in the original space, we also have $f(\mathcal{R}^{1}_{+})\leq f(\mathcal{R}^{2}_{+})$.